### v1

# On the maximum ratio between chromatic number and clique number

Abstract.
Let $f(n)$ be the maximum, over all graphs $G$ on $n$ vertices, of the ratio $\frac{\chi(G)}{\omega(G)}$, where $\chi(G)$ denotes the chromatic number of $G$ and $\omega(G)$ the clique number of $G$. In 1967, Erdős showed that
$\Big(\frac{1}{4}+o(1)\Big)\frac{n}{(\log_{2}n)^{2}}\leqslant f(n)\leqslant\big(4+o(1)\big)\frac{n}{(\log_{2}n)^{2}}.$
We show that
$f(n)\leqslant\big(c+o(1)\big)\frac{n}{(\log_{2}n)^{2}}$
for some $c<3.72$. This follows from recent improvements in the asymptotics of Ramsey numbers and is the first improvement in the asymptotics of $f(n)$ established by Erdős.
During this work, Igor Araujo was partially supported by Parker Memorial Fellowship, Schark Fellowship, and NSF RTG DMS-1937241 and Rafael Filipe was supported by CNPq.
1. Introduction
Ramsey Theory, often described as the investigation of when order arises from chaos, is one of the most influential areas in Graph Theory. The Ramsey number $R(s,t)$ is the minimum $n$ such that every red/blue edge-coloring of the complete graph on $n$ vertices contains either a clique on $s$ vertices colored red or a clique on $t$ vertices colored blue.
The first explicit upper bound for Ramsey numbers was obtained by Erdős and Szekeres [E-Sz] in 1935, who showed that $R(s,t)\leqslant\binom{s+t-2}{s-1}$. Note that in the diagonal case, this shows that $R(k,k)\leqslant 4^{k+o(k)}$. On the other hand, Erdős [Ramsey.lower] established the lower bound $R(k,k)\geqslant 2^{k/2+o(k)}$ in 1947 in one of the first uses of the probabilistic method. Subsequent improvements on the lower bound have only refined it by a constant multiplicative factor (see, e.g., Spencer’s bound obtained via the Lovász Local Lemma [Spencer.LLL]). Since then, one of the major problems in modern combinatorics has been to determine the correct asymptotic behavior of the diagonal Ramsey number $R(k,k)$. More specifically, one could ask whether the limit
$\lim_{k\to\infty}\frac{\log(R(k,k))}{k}$ (1)
exists and, if so, what is its value. This question appears as Problem 1 in [chung], as Problem 77 in https://www.erdosproblems.com/77, and on the webpage https://mathweb.ucsd.edu/~erdosproblems/.
In 1967, Erdős [erdos] asked a seemingly unrelated question. Define $f(n)$ to be the maximum, over all graphs $G$ on $n$ vertices, of the ratio $\frac{\chi(G)}{\omega(G)}$, where $\chi(G)$ denotes the chromatic number of $G$ and $\omega(G)$ the clique number of $G$. Formally, we have
$f(n)\coloneq\max\Big\{\ \frac{\chi(G)}{\omega(G)}\ :\ G\text{ is a graph on }n\text{ vertices }\Big\}.$
Erdős showed that $f(n)=\Theta(n/(\log n)^{2})$ and asked whether the following limit exists:
$\lim_{n\to\infty}\frac{f(n)}{n/(\log n)^{2}}.$ (2)
This question also appears in [erdos2], as Problem 627 in https://www.erdosproblems.com/627, as Problem 53 in [chung], and on the webpage https://mathweb.ucsd.edu/~erdosproblems/. In this note, we establish a connection between these two problems. In particular, we show that if the following conjecture is true, a solution to the former problem results in a solution to the latter problem.
Conjecture 1.1.
For every $s,t,k\in{\mathbb{N}}$ such that $st\leqslant k^{2}$ we have that
$R(s,t)\leqslant R(k,k).$
We remark that Conjecture 1.1 does not seem to appear in any existing work, but it is closely related to the Ramsey Diagonal Conjecture. In Appendix A, we briefly discuss these conjectures. The following theorem establishes a connection between the limits in (1) and (2), should they exist, under the assumption of Conjecture 1.1. Throughout this note, all logarithms are in base 2.
Theorem 1.2.
If Conjecture 1.1 holds, and $\lim\limits_{k\to\infty}\frac{\log(R(k,k))}{k}$ exists and is equal to $\mathcal{L}$, then
$f(n)=\big(\mathcal{L}^{2}+o(1)\big)\frac{n}{(\log n)^{2}}.$
In [erdos], Erdős mentioned that by their method, it would be easy to prove that
$\big(c+o(1)\big)\frac{n}{(\log n)^{2}}\leqslant f(n)\leqslant\big(C+o(1)\big)\frac{n}{(\log n)^{2}},$
for constants $c=\frac{1}{4}$ and $C=1$. However, this appears to be a typographical error, as a careful application of their method yields the constants $c=\frac{1}{4}$ and $C=4$. Aided by recent breakthrough developments in Ramsey theory, we can prove the following.
Theorem 1.3.
The function $f(n)$ satisfies
$f(n)\leqslant\big(3.71943+o(1)\big)\frac{n}{(\log n)^{2}}.$ (3)
Moreover, if Conjecture 1.1 holds, then
$f(n)\leqslant\big(3.70831+o(1)\big)\frac{n}{(\log n)^{2}}.$ (4)
This upper bound on $f(n)$ is the first improvement in the asymptotics of $f(n)$ since 1967.
Organization of the paper
In Section 2, we introduce some notation used throughout the paper and state bounds on $f(n)$, namely Theorems 2.1 and 2.2, establishing a connection between the asymptotic behavior of $f(n)$ and the Ramsey numbers $R(s,t)$. Theorem 1.2 is an immediate consequence of these results, whose proofs appear at the end of Section 2. In Section 3, we establish the numerical bounds of Theorem 1.3. In Appendix A, we discuss versions of the Ramsey Diagonal Conjecture and how they relate to Conjecture 1.1.
2. Proof of Theorem 1.2
In this section, we establish upper and lower bounds sufficient to prove Theorem 1.2. We write $h=o_{n}(t)$ if the functions $t,h\colon{\mathbb{N}}\to\mathbb{R}$ satisfy $\lim\limits_{n\to\infty}\frac{h(n)}{t(n)}=0$. When the variable $n$ is clear from the context, we simply write $h=o(t)$. Let us define
$g(n)\coloneqq\frac{(\log n)^{2}}{n}f(n),\quad L\coloneqq\liminf\limits_{k\to\infty}\;\frac{\log(R(k,k))}{k}\quad\text{ and }\quad M\coloneq\limsup_{t\to\infty}\;\max_{s\leqslant t}\ \frac{\log(R(s,t))}{\sqrt{st\ }}$
The next result establishes a lower bound for the limit inferior $\liminf\limits_{n\to\infty}g(n)$ in terms of $L$, recovering Erdős’ original lower bound $g(n)\geqslant 1/4+o(1)$, and a lower bound for the limit superior $\limsup\limits_{n\to\infty}g(n)$ in terms of $M$.
Theorem 2.1.
The function $g(n)$ satisfies
$\liminf_{n\to\infty}g(n)\geqslant L^{2}\quad\text{ and }\quad\limsup_{n\to\infty}g(n)\geqslant M^{2}.$
The following result establishes upper bounds for the limit superior $\limsup\limits_{n\to\infty}g(n)$.
Theorem 2.2.
The function $g(n)$ satisfies
$\limsup_{n\to\infty}g(n)\leqslant M^{2}.$
Before proceeding with the proofs of these theorems, we observe that
$\limsup_{k\to\infty}\;\frac{\log(R(k,k))}{k}\leqslant\limsup_{t\to\infty}\;\max_{s\leqslant t}\ \frac{\log(R(s,t))}{\sqrt{st\ }}=M.$ (5)
Fact 2.3.
Under the assumption of Conjecture 1.1, equality holds in (5).
Proof.
For $0<s\leqslant t$, let $k$ be the positive integer such that $(k-1)^{2}<st\leqslant k^{2}$. Let
$D\coloneq\limsup_{k\to\infty}\;\frac{\log(R(k,k))}{k}.$
Conjecture 1.1 implies that
$R(s,t)\leqslant R(k,k)\leqslant 2^{(D+o_{k}(1))k}$
and, therefore, we have
$\frac{\log(R(s,t))}{\sqrt{st\ }}\leqslant\frac{\log(R(k,k))}{k-1}\leqslant D+o_{k}(1)=D+o_{t}(1).\qed$
Combining Theorems 2.1 and 2.2, Fact 2.3 proves Theorem 1.2. We now proceed to prove Theorems 2.1 and 2.2, starting with Theorem 2.1.
Proof of Theorem 2.1.
For each $n\in{\mathbb{N}}$, let $k_{n}\in{\mathbb{N}}$ be such that $R(k_{n},k_{n})\leqslant n<R(k_{n}+1,k_{n}+1)$. By the definition of $L$, observe that
$\log n\geqslant(L+o_{n}(1))k_{n}.$ (6)
Let $G$ be a graph on $n$ vertices with no clique or independent set of size $k_{n}+1$. In other words, $\alpha(G)\leqslant k_{n}$ and $\omega(G)\leqslant k_{n}$. As $\chi(G)\geqslant n/\alpha(G)$, we conclude that
$g(n)=\frac{f(n)}{n/(\log n)^{2}}\geqslant\frac{\chi(G)}{\omega(G)}\cdot\frac{(\log n)^{2}}{n}\geqslant\frac{(\log n)^{2}}{\alpha(G)\omega(G)}\geqslant\frac{(\log n)^{2}}{k_{n}^{2}}\geqslant L^{2}+o_{n}(1),$ (7)
where the last inequality is true by (6). Hence, $\liminf\limits_{n\to\infty}g(n)\geqslant L^{2}$.
Moreover, we know that there exist an increasing sequence $\{t_{i}\}_{i\in{\mathbb{N}}}$ and a sequence $\{s_{i}\}_{i\in{\mathbb{N}}}$ such that $s_{i}\leqslant t_{i}$ for every $i\in{\mathbb{N}}$, and
$\lim\limits_{i\to\infty}\frac{\log(R(s_{i},t_{i}))}{\sqrt{s_{i}t_{i}\ }}=M.$
Then, for $n_{i}\coloneqq R(s_{i}+1,t_{i}+1)-1$, we have $(\log n_{i})^{2}\geqslant s_{i}t_{i}(M^{2}+o_{i}(1))$, since $n_{i}>R(s_{i},t_{i})$. Thus, a similar calculation as in (7) implies that
$g(n_{i})=\frac{f(n_{i})}{n_{i}/(\log n_{i})^{2}}\geqslant M^{2}+o_{i}(1),$
which implies that $\limsup\limits_{n\to\infty}g(n)\geqslant M^{2}$. ∎
Now we focus on proving Theorem 2.2. The proof relies on the following bound on the chromatic number of a graph due to Erdős [erdos], which is obtained by greedily picking maximum independent sets as color classes until few vertices remain, with each one receiving a new color.
Lemma 2.4.
If for every subgraph $G^{\prime}$ of $G$ on at least $m_{0}$ vertices we have that $\alpha(G^{\prime})\geqslant r$, then
$\chi(G)\leqslant\frac{n}{r}+m_{0}.$
With this lemma in hand, the proof is straightforward.
Proof of Theorem 2.2.
Let $G$ be a graph on $n$ vertices and $m\in\mathbb{N}$ be such that
$n\geqslant m\geqslant\frac{n}{(\log n)^{3}}.$
For $1\leqslant s\leqslant t$ such that $m<R(s+1,t+1)$, we have that either $s<\log t$ and
$R(s+1,t+1)\leqslant\binom{s+t}{s}=2^{o_{t}\left(\sqrt{st}\right)},$
or $s\geqslant\log t$ and
$R(s+1,t+1)\leqslant 2^{(M+o_{t}(1))\sqrt{st}},$
where the last inequality follows from the definition of $M$ and the fact that $s$ goes to infinity with $t$. Thus, in any case, we obtain
$\log m\leqslant\big(M+o_{m}(1)\big)\sqrt{st}.$ (8)
Notice that any graph $H$ on $m$ vertices provides a red/blue edge-coloring of $K_{m}$ implying that $m<R(\omega(H)+1,\alpha(H)+1)$. Thus, by (8), for every subgraph $G^{\prime}$ of $G$ on $m$ vertices, we have
$\omega(G^{\prime})\cdot\alpha(G^{\prime})\geqslant\Big(\frac{1}{M^{2}}+o_{m}(1)\Big)(\log m)^{2}.$
Since $\omega(G^{\prime})\leqslant\omega(G)$, we obtain that
$\alpha(G^{\prime})\geqslant\Big(\frac{1}{M^{2}}+o_{m}(1)\Big)\frac{(\log m)^{2}}{\omega(G)}.$
Hence, together with the fact that $\log m=\big(1+o_{n}(1)\big)\log n$, Lemma 2.4 yields
$\chi(G)\leqslant\big(M^{2}+o_{n}(1)\big)\frac{n\cdot\omega(G)}{(\log n)^{2}}+\frac{n}{(\log n)^{3}}.$
Finally, we conclude that
$\frac{\chi(G)}{\omega(G)}\leqslant\big(M^{2}+o_{n}(1)\big)\frac{n}{(\log n)^{2}}.\qed$
3. Proof of Theorem 1.3
In this section, we prove Theorem 1.3. The currently best known upper bound for $R(k,k)$ is
$R(k,k)\leqslant\left(4e^{-0.14e^{-1}}\right)^{k+o(k)}.$ (9)
This bound follows from the recent breakthrough result by Campos, Griffiths, Morris, and Sahasrabudhe [CGMS], along with its improvement by Gupta, Ndiaye, Norin, and Wei [GNNW]. Precisely, they showed that
$R(s,t)\leqslant e^{-\delta s+o(t)}\binom{s+t}{s}$ (10)
for $s\leqslant t$, where $\delta=0.14e^{-1}$. With this in hand, we can now prove Theorem 1.3.
Proof of Theorem 1.3.
If Conjecture 1.1 holds, then Theorem 2.2, Fact 2.3, and (9) imply (4). Indeed,
$f(n)\leqslant\big((\log 4e^{-0.14e^{-1}})^{2}+o(1)\big)\frac{n}{(\log n)^{2}}<\big(3.70831+o(1)\big)\frac{n}{(\log n)^{2}}.$
Thus, we now focus on proving (3). For $s\leqslant t$, let $x=\frac{s}{s+t}$. First, note that $\log R(s,t)=o(\sqrt{st})$ when $s=o(t)$. Then, we assume that $x\in[\varepsilon,1/2]$ for some $\varepsilon>0$. Observe that (10) implies
$\frac{\log(R(s,t))}{\sqrt{st\ }}\leqslant\frac{\log\binom{s+t}{s}-\delta s\log e+o_{t}(t)}{\sqrt{st}}=\frac{\log\binom{s/x}{s}-\delta s\log e}{s\sqrt{\frac{1-x}{x}}}+o_{t}(1).$
Since $\log\binom{n}{\alpha n}\leqslant H(\alpha)\cdot n$ for $\alpha\in(0,1)$, where $H(x)=-x\log x-(1-x)\log(1-x)$ is the binary entropy function, we obtain
$\frac{\log\binom{s/x}{s}-\delta s\log e}{s\sqrt{\frac{1-x}{x}}}\leqslant\frac{H(x)\cdot\frac{s}{x}-\delta s\log e}{s\sqrt{\frac{1-x}{x}}}=\frac{H(x)-\delta x\log e}{\sqrt{x(1-x)}}.$
Therefore, we have that
$\frac{\log(R(s,t))}{\sqrt{st\ }}\leqslant\frac{H(x)-\delta x\log e}{\sqrt{x(1-x)}}+o_{t}(1).$
Define $\varphi(x)\coloneq\frac{H(x)-\delta x\log e}{\sqrt{x(1-x)}}$. By differentiating, the maximum of $\varphi(x)$ is achieved for some $x\in(0,1/2]$ satisfying
$(1-x)\log(1-x)-x\log(x)-\delta x\log e=0.$
For such $x$, since $\delta=0.14e^{-1}$, we have that $\varphi(x)^{2}<3.71943$ and, by Theorem 2.2, we conclude
$\limsup_{n\to\infty}g(n)\leqslant M^{2}\leqslant\max_{x\in(0,1/2]}\varphi(x)^{2}<3.71943.\qed$
Appendix A Ramsey Diagonal Conjecture
One natural question is to understand the growth of the Ramsey Numbers $R(s,t)$ for fixed $s+t$. More precisely, one could ask whether the values of the Ramsey numbers decrease as we move away from the diagonal. A similar question can be asked in the multicolored case. The Diagonal Conjecture poses a positive answer to this question. For two colors, we summarize this conjecture in the following statement.
Conjecture A.1 (Ramsey Diagonal Conjecture (RDC)).
For every $s_{1}\leqslant s_{2}\leqslant t_{2}\leqslant t_{1}$ such that $s_{1}+t_{1}\leqslant s_{2}+t_{2}$ we have that
$R(s_{1},t_{1})\leqslant R(s_{2},t_{2}).$
It is widely believed that the Diagonal Conjecture is very difficult to prove. Even in the case $R(t-1,t+1)\leqslant R(t,t)$, there is no relevant progress. Liang, Radziszowski, and Xu [LIANG2019195] discussed several consequences of this conjecture and provided some evidence for its validity.
In [erdos], Erdős used the fact that, given some positive integer $n$, if $\binom{s+t}{t}\geqslant n$, then $st$ is minimized when $st=\lfloor{k/2}\rfloor\lfloor{(k+1)/2}\rfloor$, where $k$ is the smallest integer for which $\binom{k}{\lfloor{k/2}\rfloor}\geqslant n$. The binomial coefficient naturally arises from the first bound on the Ramsey numbers given by Erdős and Szekeres [E-Sz]. This motivates Conjecture 1.1 as a natural bridge to extend Erdős’ proof. We note that Conjecture 1.1 can be stated as a weak version of the following conjecture.
Conjecture A.2 (Multiplicative form of RDC).
For every $s_{1}\leqslant s_{2}\leqslant t_{2}\leqslant t_{1}$ such that $s_{1}t_{1}\leqslant s_{2}t_{2}$ we have that
$R(s_{1},t_{1})\leqslant R(s_{2},t_{2}).$
It is easy to see that Conjecture A.2 implies Conjecture A.1. This follows from the fact that if $s_{1}\leqslant s_{2}\leqslant t_{2}\leqslant t_{1}$ and $s_{1}t_{1}\leqslant s_{2}t_{2}$, then $s_{1}+t_{1}\leqslant s_{2}+t_{2}$.