### v2

# A note on the maximum ratio between chromatic number and clique number

Abstract.
Let $f(n)$ be the maximum, over all graphs $G$ on $n$ vertices, of the ratio $\frac{\chi(G)}{\omega(G)}$, where $\chi(G)$ denotes the chromatic number of $G$ and $\omega(G)$ the clique number of $G$. In 1967, Erdős showed that
$\Big(\frac{1}{4}+o(1)\Big)\frac{n}{(\log_{2}n)^{2}}\leqslant f(n)\leqslant\big(4+o(1)\big)\frac{n}{(\log_{2}n)^{2}}.$
We show that
$f(n)\leqslant\big(c+o(1)\big)\frac{n}{(\log_{2}n)^{2}}$
for some $c<3.72$. This follows from recent improvements in the asymptotics of Ramsey numbers and is the first improvement in the asymptotics of $f(n)$ established by Erdős.
During this work, Igor Araujo was partially supported by Parker Memorial Fellowship, Schark Fellowship, and NSF RTG DMS-1937241 and Rafael Filipe was supported by CNPq.
1. Introduction
Ramsey Theory is a fundamental area of Graph Theory concerned with the existence of unavoidable structures in large graphs. The Ramsey number $R(s,t)$ is the minimum $n$ such that every red/blue edge-coloring of the complete graph on $n$ vertices contains either a clique on $s$ vertices colored red or a clique on $t$ vertices colored blue.
While Ramsey [9] proved in 1930 that Ramsey numbers are finite, the first explicit upper bound was obtained by Erdős and Szekeres [3] in 1935, who showed that
$R(s,t)\leqslant\binom{s+t-2}{s-1}$. (1)
Note that in the diagonal case, this shows that $R(k,k)\leqslant 4^{k+o(k)}$. On the other hand, Erdős [4] established the lower bound $R(k,k)\geqslant 2^{k/2+o(k)}$ in 1947 in one of the first uses of the probabilistic method. Subsequent improvements on the lower bound have only refined it by a constant multiplicative factor (see, e.g., Spencer’s bound obtained via the Lovász Local Lemma [10]). Since then, one of the major problems in modern combinatorics has been to determine the correct asymptotic behavior of the diagonal Ramsey number $R(k,k)$. More specifically, one could ask whether the limit
$\lim_{k\to\infty}\frac{\log(R(k,k))}{k}$ (2)
exists and, if so, what is its value [^1] This question appears as Problem 1 in [2], as Problem 77 in https://www.erdosproblems.com/77, and on the webpage https://mathweb.ucsd.edu/~erdosproblems/..
In 1967, Erdős [5] asked a seemingly unrelated question. Define $f(n)$ to be the maximum, over all graphs $G$ on $n$ vertices, of the ratio $\frac{\chi(G)}{\omega(G)}$, where $\chi(G)$ denotes the chromatic number of $G$ and $\omega(G)$ the clique number of $G$. Formally, we have
$f(n)\coloneq\max\Big\{\ \frac{\chi(G)}{\omega(G)}\ :\ G\text{ is a graph on }n\text{ vertices }\Big\}.$
Erdős showed that $f(n)=\Theta(n/(\log n)^{2})$ and asked [^2] This question also appears in [6], as Problem 627 in https://www.erdosproblems.com/627, as Problem 53 in [2], and on the webpage https://mathweb.ucsd.edu/~erdosproblems/. whether the following limit exists:
$\lim_{n\to\infty}\frac{f(n)}{n/(\log n)^{2}}.$ (3)
In this note, we establish a connection between these two problems. The upper bound on $f(n)$ from [5] combines (1) with the observation that, if $\binom{s+t}{t}\geqslant n$, then
$st\geqslant\lfloor{k/2}\rfloor\lfloor{(k+1)/2}\rfloor,$
where $k$ is the smallest integer satisfying $\binom{k}{\lfloor{k/2}\rfloor}\geqslant n$. This motivates Conjecture 1.1 below as a natural bridge to extend Erdős’ proof.
Conjecture 1.1.
For every $s,t,k\in{\mathbb{N}}$ such that $st\leqslant k^{2}$ we have that
$R(s,t)\leqslant R(k,k).$
We were unable to find a previous statement of Conjecture 1.1 in the literature, but it is closely related to the Diagonal Conjecture (Conjecture 1.4; see Section 1.1 for a more detailed discussion). The following theorem establishes a connection between the limits in (2) and (3), should they exist, under the assumption of Conjecture 1.1. Throughout this note, all logarithms are in base 2.
Theorem 1.2.
If Conjecture 1.1 holds, and $\lim\limits_{k\to\infty}\frac{\log(R(k,k))}{k}$ exists and is equal to $\ell$, then
$f(n)=\big(\ell^{2}+o(1)\big)\frac{n}{(\log n)^{2}}.$
In [5], Erdős mentioned that by their method, it would be easy to prove that
$\big(c+o(1)\big)\frac{n}{(\log n)^{2}}\leqslant f(n)\leqslant\big(C+o(1)\big)\frac{n}{(\log n)^{2}},$
for constants $c=\frac{1}{4}$ and $C=1$. However, this appears to be a typographical error [^3] Indeed, if $f(n)\leqslant(1+o(1))\frac{n}{(\log n)^{2}}$ is true, then Theorem 2.1 would imply that $R(k,k)\leqslant 2^{k+o(k)}$., as a careful application of their method yields the constants $c=\frac{1}{4}$ and $C=4$.
Aided by recent breakthrough developments in Ramsey theory, we can prove the following.
Theorem 1.3.
The function $f(n)$ satisfies
$f(n)\leqslant\big(3.71943+o(1)\big)\frac{n}{(\log n)^{2}}.$ (4)
Moreover, if Conjecture 1.1 holds, then
$f(n)\leqslant\big(3.70831+o(1)\big)\frac{n}{(\log n)^{2}}.$ (5)
This upper bound on $f(n)$ is the first improvement in the asymptotics of $f(n)$ since 1967.
1.1. Conjecture 1.1 and the Diagonal Conjecture
We note that Conjecture 1.1 is closely related to the following conjecture, which is named the Diagonal Conjecture in [8].
Conjecture 1.4 (Diagonal Conjecture (DC)).
For every $s_{1}\leqslant s_{2}\leqslant t_{2}\leqslant t_{1}$ such that $s_{1}+t_{1}\leqslant s_{2}+t_{2}$ we have that
$R(s_{1},t_{1})\leqslant R(s_{2},t_{2}).$
It is widely believed that the Diagonal Conjecture is very difficult to prove. Even in the case $R(t-1,t+1)\leqslant R(t,t)$, there is no relevant progress. Observe that Conjecture 1.1 can be stated as a weak version of the following conjecture.
Conjecture 1.5 (Multiplicative form of DC).
For every $s_{1}\leqslant s_{2}\leqslant t_{2}\leqslant t_{1}$ such that $s_{1}t_{1}\leqslant s_{2}t_{2}$ we have that
$R(s_{1},t_{1})\leqslant R(s_{2},t_{2}).$
It is easy to see that if $s_{1}\leqslant s_{2}\leqslant t_{2}\leqslant t_{1}$ and $s_{1}t_{1}\leqslant s_{2}t_{2}$, then $s_{1}+t_{1}\leqslant s_{2}+t_{2}$ and therefore Conjecture 1.5 implies Conjecture 1.4.
Organization of the paper
In Section 2, we introduce some notation used throughout the paper and state bounds on $f(n)$, namely Theorems 2.1 and 2.2, establishing a connection between the asymptotic behavior of $f(n)$ and the Ramsey numbers $R(s,t)$. The proof of Theorem 1.2, which is an immediate consequence of these results, is also given in Section 2. In Section 3, we establish the numerical bounds of Theorem 1.3.
2. Proof of Theorem 1.2
In this section, we establish bounds sufficient to prove Theorem 1.2. Let us define
$g(n)\coloneqq\frac{(\log n)^{2}}{n}f(n),\quad M\coloneq\limsup_{t\to\infty}\;\max_{s\leqslant t}\ \frac{\log(R(s,t))}{\sqrt{st\ }},$
$L\coloneqq\liminf\limits_{k\to\infty}\;\frac{\log(R(k,k))}{k}\quad\text{ and }\quad D\coloneq\limsup_{k\to\infty}\;\frac{\log(R(k,k))}{k}.$
The following theorems establish more precisely the connection between limits (2) and (3). The first result gives the exactly value of $\limsup g(n)$ in terms of the Ramsey numbers.
Theorem 2.1.
The function $g(n)$ satisfies
$\limsup_{n\to\infty}g(n)=M^{2}.$
The second one establishes a lower bound on $\liminf g(n)$ in terms of the Ramsey numbers.
Theorem 2.2.
The function $g(n)$ satisfies
$\liminf_{n\to\infty}g(n)\geqslant L^{2}.$
Before proceeding to the proof of these results, let us show how to derive Theorem 1.2.
Proof of Theorem 1.2.
The assumption that $\lim\frac{\log R(k,k)}{k}$ exists and equals $\ell$ implies $L=D=\ell$. Thus, by Theorems 2.1 and 2.2, it suffices to prove that $D=M$. To this end, observe that restricting the maximum to $s=t$ in the definition of $M$ yields
$D=\limsup_{k\to\infty}\;\frac{\log(R(k,k))}{k}\leqslant\limsup_{t\to\infty}\;\max_{s\leqslant t}\ \frac{\log(R(s,t))}{\sqrt{st\ }}=M.$ (6)
On the other hand, if $0<s\leqslant t$ and $k$ is the positive integer such that $(k-1)^{2}<st\leqslant k^{2}$, Conjecture 1.1 implies that
$R(s,t)\leqslant R(k,k)\leqslant 2^{(D+o_{k}(1))k}.$
Therefore, we have
$\frac{\log(R(s,t))}{\sqrt{st\ }}\leqslant\frac{\log(R(k,k))}{k-1}\leqslant D+o_{k}(1)=D+o_{t}(1),$
which implies $M\leqslant D$. Combined with (6), we conclude that $D=M$. ∎
We now proceed to prove Theorems 2.1 and 2.2, starting with Theorem 2.2.
Proof of Theorem 2.2.
For each $n\in{\mathbb{N}}$, let $k_{n}\in{\mathbb{N}}$ be such that $R(k_{n},k_{n})\leqslant n<R(k_{n}+1,k_{n}+1)$. By the definition of $L$, observe that
$\log n\geqslant\big(L+o_{n}(1)\big)k_{n}.$ (7)
Let $G$ be a graph on $n$ vertices with no clique or independent set of size $k_{n}+1$. In other words, $\alpha(G)\leqslant k_{n}$ and $\omega(G)\leqslant k_{n}$. As $\chi(G)\geqslant n/\alpha(G)$, we conclude that
$g(n)=\frac{f(n)}{n/(\log n)^{2}}\geqslant\frac{\chi(G)}{\omega(G)}\cdot\frac{(\log n)^{2}}{n}\geqslant\frac{(\log n)^{2}}{\alpha(G)\omega(G)}\geqslant\frac{(\log n)^{2}}{k_{n}^{2}}\geqslant L^{2}+o_{n}(1),$
where the last inequality is true by (7). Hence, $\liminf g(n)\geqslant L^{2}$. ∎
Now we focus on proving Theorem 2.1. The proof relies on the following bound on the chromatic number of a graph from [5], which is obtained by greedily picking maximum independent sets as color classes until few vertices remain, with each remaining vertex receiving a new color.
Lemma 2.3.
If $\alpha(G^{\prime})\geqslant r$ for every subgraph $G^{\prime}\subset G$ with at least $m_{0}$ vertices, then
$\chi(G)\leqslant\frac{n}{r}+m_{0}.$
With this lemma in hand, the proof is straightforward.
Proof of Theorem 2.1.
Let $G$ be a graph on $n$ vertices and $m\in\mathbb{N}$ be such that
$n\geqslant m\geqslant\frac{n}{(\log n)^{3}}.$
For $1\leqslant s\leqslant t$ such that $m<R(s+1,t+1)$, we have that either $s<\log t$ and
$R(s+1,t+1)\leqslant\binom{s+t}{s}=2^{o_{t}\left(\sqrt{st}\right)},$
or $s\geqslant\log t$ and
$R(s+1,t+1)\leqslant 2^{(M+o_{t}(1))\sqrt{st}},$
where the last inequality follows from the definition of $M$ and the fact that $s$ goes to infinity with $t$. Thus, in any case, we obtain
$\log m\leqslant\big(M+o_{m}(1)\big)\sqrt{st}.$ (8)
Notice that any graph $H$ on $m$ vertices provides a red/blue edge-coloring of $K_{m}$ implying that $m<R(\omega(H)+1,\alpha(H)+1)$. Thus, by (8), for every subgraph $G^{\prime}$ of $G$ on $m$ vertices, we have
$\omega(G^{\prime})\cdot\alpha(G^{\prime})\geqslant\Big(\frac{1}{M^{2}}+o_{m}(1)\Big)(\log m)^{2}.$
Since $\omega(G^{\prime})\leqslant\omega(G)$, we obtain that
$\alpha(G^{\prime})\geqslant\Big(\frac{1}{M^{2}}+o_{m}(1)\Big)\frac{(\log m)^{2}}{\omega(G)}.$
Hence, together with the fact that $\log m=\big(1+o_{n}(1)\big)\log n$, Lemma 2.3 yields
$\chi(G)\leqslant\big(M^{2}+o_{n}(1)\big)\frac{n\cdot\omega(G)}{(\log n)^{2}}+\frac{n}{(\log n)^{3}}.$
Then, we conclude that
$\frac{\chi(G)}{\omega(G)}\leqslant\big(M^{2}+o_{n}(1)\big)\frac{n}{(\log n)^{2}}.$ (9)
Moreover, we know that there exist an increasing sequence $\{t_{i}\}_{i\in{\mathbb{N}}}$ and a sequence $\{s_{i}\}_{i\in{\mathbb{N}}}$ such that $s_{i}\leqslant t_{i}$ for every $i\in{\mathbb{N}}$, and
$\lim\limits_{i\to\infty}\frac{\log(R(s_{i},t_{i}))}{\sqrt{s_{i}t_{i}\ }}=M.$
Then, for $n_{i}\coloneqq R(s_{i}+1,t_{i}+1)-1$, we have $(\log n_{i})^{2}\geqslant s_{i}t_{i}\big(M^{2}+o_{i}(1)\big)$, since $n_{i}>R(s_{i},t_{i})$. The definition of $n_{i}$ also implies that there is a graph $G_{i}$ on $n_{i}$ vertices satisfying $\alpha(G_{i})\leqslant s_{i}$ and $\omega(G_{i})\leqslant t_{i}$. Thus we have
$g(n_{i})=\frac{f(n_{i})}{n_{i}/(\log n_{i})^{2}}\geqslant\frac{\chi(G_{i})}{\omega(G_{i})}\cdot\frac{(\log n_{i})^{2}}{n_{i}}\geqslant\frac{(\log n_{i})^{2}}{\alpha(G_{i})\omega(G_{i})}\geqslant\frac{(\log n_{i})^{2}}{s_{i}t_{i}}\geqslant M^{2}+o_{i}(1),$
which together with (9) implies that $\limsup g(n)=M^{2}$.∎
3. Proof of Theorem 1.3
In this section, we prove Theorem 1.3. The currently best known upper bound for $R(k,k)$ is
$R(k,k)\leqslant\big(4e^{-0.14e^{-1}}\big)^{k+o(k)}.$ (10)
This bound follows from the recent breakthrough result by Campos, Griffiths, Morris, and Sahasrabudhe [1], along with its improvement by Gupta, Ndiaye, Norin, and Wei [7]. Precisely, they showed that
$R(s,t)\leqslant e^{-\delta s+o(t)}\binom{s+t}{s}$ (11)
for $s\leqslant t$, where $\delta=0.14e^{-1}$. With this in hand, we can now prove Theorem 1.3.
Proof of Theorem 1.3.
As seen in the proof of Theorem 1.2, if Conjecture 1.1 holds, then $D=M$. Hence, Theorem 2.1 and (10) imply (5). Indeed,
$f(n)\leqslant\big((\log 4e^{-0.14e^{-1}})^{2}+o(1)\big)\frac{n}{(\log n)^{2}}<\big(3.70831+o(1)\big)\frac{n}{(\log n)^{2}}.$
Thus, we now focus on proving (4). For $s\leqslant t$, let $x=\frac{s}{s+t}$. First, note that $\log R(s,t)=o(\sqrt{st})$ when $s=o(t)$. Then, we assume that $x\in[\varepsilon,1/2]$ for some $\varepsilon>0$. Observe that (11) implies
$\frac{\log(R(s,t))}{\sqrt{st\ }}\leqslant\frac{\log\binom{s+t}{s}-\delta s\log e+o_{t}(t)}{\sqrt{st}}=\frac{\log\binom{s/x}{s}-\delta s\log e}{s\sqrt{\frac{1-x}{x}}}+o_{t}(1).$
Since $\log\binom{n}{\alpha n}\leqslant H(\alpha)\cdot n$ for $\alpha\in(0,1)$, where $H(x)=-x\log x-(1-x)\log(1-x)$ is the binary entropy function, we obtain
$\frac{\log\binom{s/x}{s}-\delta s\log e}{s\sqrt{\frac{1-x}{x}}}\leqslant\frac{H(x)\cdot\frac{s}{x}-\delta s\log e}{s\sqrt{\frac{1-x}{x}}}=\frac{H(x)-\delta x\log e}{\sqrt{x(1-x)}}.$
Therefore, we have that
$\frac{\log(R(s,t))}{\sqrt{st\ }}\leqslant\frac{H(x)-\delta x\log e}{\sqrt{x(1-x)}}+o_{t}(1).$
Define $\varphi(x)\coloneq\frac{H(x)-\delta x\log e}{\sqrt{x(1-x)}}$. By differentiating, the maximum of $\varphi(x)$ is achieved for some $x\in(0,1/2]$ satisfying
$(1-x)\log(1-x)-x\log(x)-\delta x\log e=0.$
For such $x$, since $\delta=0.14e^{-1}$, we have that $\varphi(x)^{2}<3.71943$ and, by Theorem 2.1, we conclude
$\limsup_{n\to\infty}g(n)=M^{2}\leqslant\max_{x\in(0,1/2]}\varphi(x)^{2}<3.71943.\qed$
Acknowledgements
The authors thank Rob Morris for his helpful suggestions that heavily improved the presentation of this paper.